# Consistent Validation for Predictive Methods in Spatial Settings

### v1

@@ B.3.4 Restatement and Proof of Inconsistency of $1$-nearest neighbor
$\mathbb{E}|R^{\chi}_{Q^{\textup{{test}}}}(h)-(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|^{2}\geq\mathbb{E}|(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)-\mathbb{E}[(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)]|^{2}.$ (46)
Because the $\epsilon^{\textup{val}}_{n}$ are independent the variance of $(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}$ is additive
$\mathbb{E}|(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)-\mathbb{E}(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|^{2}=\sum_{n=1}^{N^{\textup{{val}}}}(w^{\textup{{NN}},\!1}_{n})^{2}\mathbb{E}[\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n}))-e^{\chi}_{h}(S^{\textup{{val}}}_{n})]\geq V\sum_{n=1}^{N^{\textup{{val}}}}(w^{\textup{{NN}},\!1}_{n})^{2}.$ (47)
where $40<V<\Delta^{2}/$ 4 is the assumed lower bound on the variance of $\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n}))-e^{\chi}_{h}(S^{\textup{{val}}}_{n})$ and we have left implicit the dependence of the weights on $N^{\textup{{val}}}$. Also, |$|\{1\leq n\leq N^{\textup{{val}}}:d_{\mathcal{S}}(S^{\prime},S^{\textup{{val}}}_{n})=r|\leq 1$, implies $S^{\prime}$ has exactly one $1$-nearest neighbor in $S^{\textup{{val}}}_{1:N^{\textup{{val}}}}$, call the index of this neighbor $n(S^{\prime})$. Then
Combining Eqn. 47 and Eqn. 48
$\mathbb{E}[|R^{\chi}_{Q^{\textup{{test}}}}(h)-(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|^{2}]\geq VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}.$ (49)
We now apply Prop. B.5 with $U=|R^{\chi}_{Q^{\textup{{test}}}}(h)-(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|$ to conclude that for $\delta\in(1-\frac{VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}}{\Delta^{2}},1)$ with probability at least $1-\delta$
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|\geq\Delta\sqrt{1-\frac{1-\frac{VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}}{\Delta^{2}}}{\delta}}>0.$ (50)
As neither $\delta$ nor the right hand side of Eqn. 50 depend on $N^{\textup{{val}}}$, $1$-nearest neighbor is not consistent under infill asymptotics. ∎
B.4 General Nearest Neighbor Bound and Selecting the Number of Neighbors
@@
In this section, we prove two results giving bounds on the performance of nearest neigbhor risk estimation that will be the basis of later results. The first, already stated in the main text, is a general bound for any $k$.
@@
The second result we will show is specific to SNN and relates the error incurred by using $k^{\star}_{T_{2}}$ to the error of the minimizer of the bound from Thm. 5.1:
nnminimizationpowtwo
B.4.1 Preliminary Result: Hoeffding’s Inequality
@@
Lemma B.6 (Hoeffding’s Inequality, (Hoeffding, 1963, Theorem 2)).
@@
B.4.2 Proof of General Nearest Neighbor Risk Estimation Bound
We again recall and prove Thm. 5.1. The idea is to apply triangle inequality to split the error into a bias term and a sampling error term. The sampling error term is then controlled with Lemma B.6 since the loss is bounded. The bias term is controlled using Assumption 2.5. See 5.1
@@ B.4.2 Proof of General Nearest Neighbor Risk Estimation Bound
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)|$ $\leq\underbrace{\big{|}R^{\chi}_{Q^{\textup{{test}}}}(h)-\sum_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}e^{\chi}_{h}(S^{\textup{{val}}}_{n})\big{|}}_{\tau_{1}}+\underbrace{\Big{|}\sum_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}(\ell(f(S^{\textup{{val}}}_{n},\chi(S^{\textup{{val}}}_{n}),\epsilon^{\textup{val}}_{n}),h^{\chi}(S^{\textup{{val}}}_{n}))-e^{\chi}_{h}(S^{\textup{{val}}}_{n})\Big{|}}_{\tau_{2}}.$ (53)
The first term, $\tau_{1}$ is a bias term, while the second term, $\tau_{2}$ is a sum of $N^{\textup{{val}}}$ independent variables with expectation zero. Using Assumption 2.1, we apply Hoeffding’s inequality (Lemma B.6) to bound $\tau_{2}$: for any $\delta\in(0,1)$ with probability at least $1-\delta$
@@
Recalling the definition of $w^{\textup{{NN}},k}$ (Def. 4.2) and that each $A^{k}(s)$ contains at least $k$ points by construction,
@@
Substituting this in Eqn. 56 and using Eqn. 55
@@
$\tau_{1}=\left\lvert\frac{1}{M^{\textup{{test}}}}\sum_{m=1}^{M^{\textup{{test}}}}(e^{\chi}_{h}(S^{\textup{{test}}}_{m})-\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{nm}^{k}e^{\chi}_{h}(S^{\textup{{val}}}_{n}))\right\rvert\leq\max_{1\leq m\leq M^{\textup{{test}}}}\left\lvert e^{\chi}_{h}(S^{\textup{{test}}}_{m})-\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{nm}^{k}e^{\chi}_{h}(S^{\textup{{val}}}_{n})\right\rvert.$ (60)
@@
$\left\lvert e^{\chi}_{h}(S^{\textup{{test}}}_{m})-\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{mn}^{k}e^{\chi}_{h}(S^{\textup{{val}}}_{n})\right\rvert=\left\lvert\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{mn}^{k}(e^{\chi}_{h}(S^{\textup{{test}}}_{m})-e^{\chi}_{h}(S^{\textup{{val}}}_{n}))\right\rvert\leq\max_{n:\alpha_{mn}^{k}>0}|e^{\chi}_{h}(S^{\textup{{test}}}_{m})-e^{\chi}_{h}(S^{\textup{{val}}}_{n})|.$ (61)
@@
The result follows from combining Eqns. 53, 54, 63 and 59. ∎
B.4.3 Proofs Related to Selecting the Number of Neighbors
We now restate and prove a bound that upper bounds the error of risk estimation with SNN (i.e. using $k^{\star}_{T_{2}}$ neighbors) to the minimum of the upper bound from Thm. 5.1 over all $k$. A key observation is that because Thm. 5.1 is conditional on the test locations, it can be minimized without the need to take a union bound over all $k$ in the set we minimize over. We first need a preliminary result, which holds for minimization over any subset of $\{1,\dots,N^{\textup{{val}}}\}$.
Proposition B.7 (Minimization of Upper Bound).
@@ B.4.3 Proofs Related to Selecting the Number of Neighbors
||| $R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T}}}(h)|\leq C_{L}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}\right)$ (64)
@@
Proof of Prop. B.7.
Because the minimization problem does not depend on a quantity that is treated as random in Thm. 5.1, we may directly apply Thm. 5.1 to $k^{\star}_{T}$ to conclude with probability at least $1-\delta$
$|R^{\chi}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T}}}(h)|\leq L\rho_{k^{\star}_{T}}+\Delta\|w^{\textup{{{NN}}},k^{\star}_{T}}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$ (66)
@@
Combining Eqns. 68, 70 and 66 gives the result. ∎
Proposition B.8 (Minimization over Powers of Two).
||| $R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq\sqrt{2}C_{L}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\right),$ (71)
where $B(S,r)$ denotes the ball of radius $r$ centered at $S$.
@@
$\min_{k\in T_{2}}\left(\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\right)$ $\leq\rho_{\tilde{k}}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{\tilde{k}})))}{\tilde{k}}}$ (75)
$\leq\rho_{k^{\star}_{T}}+\beta_{\delta}\sqrt{2}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k^{\star}_{T}}))}{k^{\star}_{T}}}$ (76)
$\leq\sqrt{2}(\rho_{k^{\star}_{T}}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k^{\star}_{T}}))}{k^{\star}_{T}}})$ (77)
$=\sqrt{2}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\right).$ (78)
The result now follows from Prop. B.7. ∎
B.5 Consistency of our Nearest Neighbor Method under Infill Asymptotics
In this section we prove Cor. 5.2, which establishes the consistency of SNN under infill asymptotics. The idea of the proof is to upper bound the $k$th order fill distance of the validation points in the test points to the fill distance of the validation points in $[0,1]^{d}$. This allows together with Thm. 5.1 and Prop. B.8 allows us to derive an upper bound on the error of our method in terms of the fill distance, from which the result follows.
B.5.1 Preliminary Lemma: Relating Fill Distances
Proposition B.9.

### v2

@@ D.3.4 Proof of Inconsistency of $1$-nearest neighbor
Because $S^{\textup{{val}}}_{m}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}\mu$, and $\mu$ has Lebesgue density, the nearest neighbor to $S$ is almost surely unique. This implies that $\hat{R}_{\textup{{NN}},1}(h)\in\{0,1\}$. Therefore,
$|\hat{R}_{\textup{{NN}},1}(h)-R_{Q^{\textup{{test}}}}(h)|\geq\min_{a\in\{0,1\}}|a-\mathbb{E}[Y^{\textrm{test}}]|=\min(\mathbb{E}[Y^{\textrm{test}}],1-\mathbb{E}[Y^{\textrm{test}}]).$ (44)
D.3.5 Inconsistency of Nearest Neighbors with Number of Neighbors Depending On Number of Validation Points
We now restate and prove Proposition 4.4, which states that nearest-neighbor risk estimation with the number of neighbors depending (only) on the number of validation points is inconsistent under infill asymptotics, regardless of type of dependence.
See 4.4
Proof.
The idea is that either 1.) $(k_{n})_{n=1}^{\infty}$ has a bounded sub-sequence, in which case along this sub-sequence, the $\hat{R}_{N^{\textup{{val}}}}(h)$ can have a variance bounded below by 00, and so by Proposition D.8 these estimators are bounded away from the test risk with fixed probability. Or 2.) the number of neighbors used tends to infinity, in which case we can find a sequence of data that accumulates more slowly around the test site, leading to many neighbors far from the point being used in the estimator, and therefore non-negligible bias.
We split into these two cases, and give an example showing in either case $\hat{R}_{N^{\textup{{val}}}}$ can be inconsistent.
Case 1: $\lim\inf_{n\to\infty}k_{n}<\infty$.
Consider a data-generating process with no covariates, $S^{\textup{{val}}}\sim U(0,1)$, $S^{\textup{{test}}}=\{\frac{1}{2}\}$, $Y_{n}|S_{n}=\epsilon_{n}\sim U(-1/2,1/2)$, $h=0$ and $\ell(y,y^{\prime})=|y-y^{\prime}|$. Because $\lim\inf_{n\to\infty}k_{n}<\infty$, there exists a $C>0$ such that $(k_{n})_{n=1}^{\infty}$ contains a bounded sub-sequence $(\tilde{k}_{n})_{n=1}^{\infty}$ with $\tilde{k}_{n}\leq C$ for all $C>0$. Since the limit superior of a sub-sequence cannot be larger than of the full sequence
$\lim\sup_{N^{\textup{{val}}}\to\infty}\mathrm{Var}(\hat{R}_{N^{\textup{{val}}}}(h))\geq\lim\sup_{n\to\infty}\mathrm{Var}(\tilde{R}_{n}(h)).$ (45)
where $\tilde{R}_{n}$ denotes the sub-sequence of $\hat{R}_{N^{\textup{{val}}}}$ where the number of validation points runs along the sub-sequence corresponding to $(\tilde{k}_{n})_{n=1}^{\infty}$.
Almost surely, for any $N^{\textup{{val}}}$ the test point $1/2$ has exactly $k_{N^{\textup{{val}}}}$-nearest neighbors, because the probability that two validation points are equidistant from $1/2$ is 00. Since a countable union of almost sure events is also an almost sure event, with probability $1$ for all $N^{\textup{{val}}}$ the test point at $1/2$ has exactly $k_{N^{\textup{{val}}}}$ neighbors. Therefore, with probability $1$, for all $N^{\textup{{val}}}$ the vector of weights $w^{\textup{{NN}},k}$ has exactly $k_{N^{\textup{{val}}}}$ non-zero entries, each with value $1/k_{N^{\textup{{val}}}}$. We condition on this probability $1$ event moving forward. In this case,
$\mathrm{Var}(\hat{R}_{N^{\textup{{val}}}})$ $=\mathrm{Var}(\sum_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}|\epsilon_{n}|)$ (46)
$=\frac{1}{k_{N^{\textup{{val}}}}^{2}}\sum_{i=1}^{k_{N^{\textup{{val}}}}}\mathrm{Var}(|\tilde{\epsilon}_{i}|)$ (47)
$=\frac{1}{48k_{N^{\textup{{val}}}}}$ (48)
where $(\tilde{\epsilon})_{i=1}^{k_{N^{\textup{{val}}}}}$ are the subset of $(\epsilon_{n})_{n=1}^{N^{\textup{{val}}}}$ corresponding to the $k_{N^{\textup{{val}}}}$ nearest neighbors to $1/2$. The factor of 48 comes from the variance of a uniform random variable on $[0,1/2]$.
For all $N^{\textup{{val}}}$ corresponding to the bounded sub-sequence $(\tilde{k}_{n})_{n=1}^{\infty}$, we conclude
$\mathrm{Var}(\tilde{R}_{n}(h))\geq\frac{1}{48C}.$ (49)
Therefore,
$\lim\sup_{N^{\textup{{val}}}\to\infty}\mathrm{Var}(\hat{R}_{N^{\textup{{val}}}}(h))\geq\frac{1}{48C}.$ (50)
Applying Proposition D.8 to the random variable $|\hat{R}_{N^{\textup{{val}}}}(h)-R(h)|$ we conclude there exists an $\epsilon,\delta$ such that with probability at least $1-\delta$,
$\lim\sup_{N^{\textup{{val}}}\to\infty}|\hat{R}_{N^{\textup{{val}}}}(h)-R(h)|\geq\epsilon.$ (51)
Case 2: $\lim\inf_{n\to\infty}k_{n}=\infty$.
Consider the data generating process, $Y_{n}|S_{n}=S_{n}$ on $[0,1]$ with $S^{\textup{{test}}}=\{0\}$ and $\ell(y,y^{\prime})=|y-y^{\prime}|$ and $h=0$. We then have $R(h)=0$. We will construct a sequence of validation sites $S^{\textup{{val}}}$ such that for each $N^{\textup{{val}}}$ less than $k_{N^{\textup{{val}}}}/2$ of the validation sites fall in the interval $[0,1/4)$. For any such sequence (supposing such a sequence exists for the moment),
$|\hat{R}_{N^{\textup{{val}}}}(h)-R(h)|=|\hat{R}_{N^{\textup{{val}}}}(h)|\geq\frac{1}{k_{N^{\textup{{val}}}}}\cdot\frac{k_{N^{\textup{{val}}}}}{2}\frac{1}{4}\geq\frac{1}{8}.$ (52)
All that remains is to construct such a sequence that also satisfies infill asymptotics. Define the function $\psi:\mathbb{N}\to(0,1)$ by
$\psi(i)=\frac{i-2^{\lfloor\log_{2}i\rfloor+1}}{2^{\lfloor\log_{2}i\rfloor+1}}.$ (53)
This corresponds to the dyadic sequence $(1/2,1/4,3/4,1/8,3/8,5/8,7/8,1/16,\dotsc)$. The essential properties of this function for our application is that the image of the function is dense on $(0,1)$ and for any $i$ at least $i/2$ of $(\phi(j))_{j=1}^{i}$ are in the interval $(0,1/2]$.
Algorithm 1 Algorithm defining $(S^{\textup{{val}}}_{n})_{n=1}^{\infty}$
  while True: do
     if $j<k_{n}$ and $n-j>j$ then
        $S^{\textup{{val}}}_{n}=\frac{1}{2}-\frac{1}{2}\psi(j)$
        $j\leftarrow j+1$.
     else
        $S^{\textup{{val}}}_{n}=\frac{1}{2}+\frac{1}{2}\phi(n-j)$
        $n\leftarrow n+1$.
     end if
  end while
The validation points are defined algorithmically via Algorithm 1. Because $k_{n}$ is unbounded, the first condition must be called infinitely often, and so $j$ eventually takes on all natural numbers in this loop.
Because $n-j>j$ each time the first condition is called, $n-j$ is incremented if and only if $j$ is not incremented $n-j$ also takes on all natural numbers in this loop. Therefore, $(S^{\textup{{val}}}_{n})_{n=1}^{\infty}$ is a dense set in $[0,1]$ because $\frac{1}{2}-\frac{1}{2}\psi(\mathbb{N})$ is dense on $[0,1/2]$ and $\frac{1}{2}+\frac{1}{2}\psi(\mathbb{N})$ is dense on $[1/2,1]$. We conclude this sequence satisfies the infill asymptotics.
For any $N^{\textup{{val}}}$, the number of validation points less than $1/2$ is not more than $k_{N^{\textup{{val}}}}$ by induction and using the first condition in the if statement. Of the points placed in $(0,1/2)$ at most half of them are in $(0,1/4]$, by our earlier observation that for any $i$ at least $i/2$ of $(\psi(j))_{j=1}^{i}$ are in the interval $(0,1/2]$. Therefore, not more than $k_{N^{\textup{{val}}}}/2$ points are in $[0,1/4)$ for any $N^{\textup{{val}}}$. ∎
D.4 General Nearest Neighbor Bound and Selecting the Number of Neighbors
@@
In this section, we prove two results giving bounds on the performance of nearest neighbor risk estimation that will be the basis of later results. The first, already stated in the main text, is a general bound for any $k$.
@@
The second result we will show is specific to SNN and relates the error incurred by using $k^{\star}_{T_{2}}$ to the error of the minimizer of the bound from Theorem 5.1:
Proposition D.10 (Minimization over Powers of Two).
||| $R_{Q^{\textup{{test}}}}(h)-\hat{R}_{\textup{{{NN}}},k^{\star}_{T_{2}}}(h)|\leq\sqrt{2}C_{L}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\right),$ (54)
where $B(S,r)$ denotes the ball of radius $r$ centered at $S$.
D.4.1 Preliminary Result: Hoeffding’s Inequality
@@
Lemma D.11 (Hoeffding’s Inequality, [Hoeffding, 1963, Theorem 2]).
@@
D.4.2 Proof of General Nearest Neighbor Risk Estimation Bound
We again recall and prove Theorem 5.1. The idea is to apply triangle inequality to split the error into a bias term and a sampling error term. The sampling error term is then controlled with Lemma D.11 since the loss is bounded. The bias term is controlled using Assumption 2.5. See 5.1
@@ D.4.2 Proof of General Nearest Neighbor Risk Estimation Bound
$|R_{Q^{\textup{{test}}}}(h)-\hat{R}_{\textup{{NN}},k}(h)|$ $\leq\underbrace{\big{|}R_{Q^{\textup{{test}}}}(h)-\sum_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}e_{h}(S^{\textup{{val}}}_{n})\big{|}}_{\tau_{1}}$ (57)
$+\underbrace{\Big{|}\sum_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}(\ell(f(S^{\textup{{val}}}_{n},\chi(S^{\textup{{val}}}_{n}),\epsilon^{\textup{val}}_{n}),h^{\chi}(S^{\textup{{val}}}_{n}))-e_{h}(S^{\textup{{val}}}_{n})\Big{|}}_{\tau_{2}}.$ (58)
The first term, $\tau_{1}$ is a bias term, while the second term, $\tau_{2}$ is a sum of $N^{\textup{{val}}}$ independent variables with expectation zero. Using Assumption 2.1, we apply Hoeffding’s inequality (Lemma D.11) to bound $\tau_{2}$: for any $\delta\in(0,1)$ with probability at least $1-\delta$
@@
Recalling the definition of $w^{\textup{{NN}},k}$ (Definition 4.2) and that each $A^{k}(s)$ contains at least $k$ points by construction,
@@
Substituting this in Equation 61 and using Equation 60
@@
$\tau_{1}$ $=\left\lvert\frac{1}{M^{\textup{{test}}}}\sum_{m=1}^{M^{\textup{{test}}}}(e_{h}(S^{\textup{{test}}}_{m})-\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{nm}^{k}e_{h}(S^{\textup{{val}}}_{n}))\right\rvert$ (65)
$\leq\max_{1\leq m\leq M^{\textup{{test}}}}\left\lvert e_{h}(S^{\textup{{test}}}_{m})-\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{nm}^{k}e_{h}(S^{\textup{{val}}}_{n})\right\rvert.$ (66)
@@
$\left\lvert e_{h}(S^{\textup{{test}}}_{m})-\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{mn}^{k}e_{h}(S^{\textup{{val}}}_{n})\right\rvert$ $=\left\lvert\sum_{n=1}^{N^{\textup{{val}}}}\alpha_{mn}^{k}(e_{h}(S^{\textup{{test}}}_{m})-e_{h}(S^{\textup{{val}}}_{n}))\right\rvert$ (67)
$\leq\max_{n:\alpha_{mn}^{k}>0}|e_{h}(S^{\textup{{test}}}_{m})-e_{h}(S^{\textup{{val}}}_{n})|.$ (68)
@@
The result follows from combining Equations 58, 59, 70 and 64. ∎
D.4.3 Proofs Related to Selecting the Number of Neighbors
We now restate and prove a bound that upper bounds the error of risk estimation with SNN (i.e. using $k^{\star}_{T_{2}}$ neighbors) to the minimum of the upper bound from Theorem 5.1 over all $k$. A key observation is that because Theorem 5.1 is conditional on the test locations, it can be minimized without the need to take a union bound over all $k$ in the set we minimize over. We first need a preliminary result, which holds for minimization over any subset of $\{1,\dots,N^{\textup{{val}}}\}$.
Proposition D.12 (Minimization of Upper Bound).
@@ D.4.3 Proofs Related to Selecting the Number of Neighbors
$|R_{Q^{\textup{{test}}}}(h)-\hat{R}_{\textup{{{NN}}},k^{\star}_{T}}(h)|$ $\leq C_{L}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}\right)$ (71)
@@
Proof of Proposition D.12.
Because the minimization problem does not depend on a quantity that is treated as random in Theorem 5.1, we may directly apply Theorem 5.1 to $k^{\star}_{T}$ to conclude with probability at least $1-\delta$
$|R(h)-\hat{R}_{\textup{{{NN}}},k^{\star}_{T}}(h)|\leq L\rho_{k^{\star}_{T}}+\Delta\|w^{\textup{{{NN}}},k^{\star}_{T}}\|_{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$ (73)
@@
Combining Equations 75, 77 and 73 gives the result. ∎
See D.10
@@
$\min_{k\in T_{2}}\Bigg{(}\rho_{k}+\beta_{\delta}$ $\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\Bigg{)}\leq\rho_{\tilde{k}}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{\tilde{k}})))}{\tilde{k}}}$ (81)
$\phantom{\qquad\qquad\qquad\qquad}\leq\rho_{k^{\star}_{T}}+\beta_{\delta}\sqrt{2}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k^{\star}_{T}}))}{k^{\star}_{T}}}$ (82)
$\phantom{\qquad\qquad\qquad\qquad}\leq\sqrt{2}(\rho_{k^{\star}_{T}}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k^{\star}_{T}}))}{k^{\star}_{T}}})$ (83)
$\phantom{\qquad\qquad\qquad\qquad}=\sqrt{2}\left(\min_{k\in T}\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\right).$ (84)
The result now follows from Proposition D.12. ∎
D.5 Consistency of our Nearest Neighbor Method under Infill Asymptotics
In this section we prove Corollary 5.2, which establishes the consistency of SNN under infill asymptotics. The idea of the proof is to upper bound the $k$th order fill distance of the validation points in the test points to the fill distance of the validation points in $[0,1]^{d}$. This allows together with Theorem 5.1 and Proposition D.10 allows us to derive an upper bound on the error of our method in terms of the fill distance, from which the result follows.
D.5.1 Preliminary Lemma: Relating Fill Distances
Proposition D.13.